# Convergence Analysis of An Alternating Nonlinear GMRES on Linear Systems

### v1

@@ 2.1.2 Symmetric and positive definite case
In the following, we consider $M$ being symmetric and positive definite (SPD), i.e., there exists a unitary matrix $U$ such that $M=U\Lambda U^{T}$, where $\Lambda$ is a diagonal matrix. Then, $e_{j+1}=M^{j+1}e_{0}$ and $e_{j+1}-e_{j}=M^{j}(M-I)e_{0}=U\Lambda^{j}(\Lambda-I)U^{T}e_{0}$.
@@
Assume that $M$ is SPD, i.e., $M=U\Lambda U^{T}$, where $U$ is a unitary matrix and $\Lambda$ is a diagonal matrix, the eigenvalues of $M$ are contained in an interval $[a,b]$ that does not include 00 and 1, and $m<k$. Then, the 2-norm of the error $e_{k+1}$ obtained from one-step NGMRES($m$) satisfies
$\|\Theta^{-1}U^{T}e_{k+1}\|\leq\epsilon(a,b,m+1)\cdot\|\Theta^{-1}U^{T}Me_{k}\|,$ (15)
@@
$\leq\min_{p_{m}\in P_{m}}\max_{t\in[a,b]}\|I-p_{m}(t)t^{-m-1}(1-t)\|\|\Theta^{-1}U^{T}Me_{k}\|$
$\leq\min_{p_{m}\in P_{m}}\max_{t\in[1/b,1/a]}\|I-p_{m}(1/t)t^{m}(t-1)\|\|\Theta^{-1}U^{T}Me_{k}\|$
$\leq\min_{p_{m}\in P_{m}}\max_{t\in[1/b,1/a]}\|I-p_{m}(t)(t-1)\|\|\Theta^{-1}U^{T}Me_{k}\|$
$\leq\min_{q_{m+1}\in P_{m+1},q_{m+1}(1)=1}\max_{t\in[1/b,1/a]}\|q_{m+1}(t)\|\|\Theta^{-1}U^{T}Me_{k}\|.$
@@
The second result can directly obtained from Theorem 2.1 using the fact that $\kappa_{2}(U)=1$. The last result, (17), is obtained from Theorem 2.2.
@@
Note that [23] studied one-step Anderson acceleration for the symmetric case. The Chebyshev polynomial $\epsilon(a,b,m+1)$ for NGMRES($m$) is degree $m+1$, but it is $m$ for AA($m$), and the $\Theta$ in [23] is $\Lambda(\Lambda-I)^{-1}$.
@@ 3 Alternating NGMRES
Note that NGMRES($m$) requires solving a least-squares problem at each iteration, which is computational expensive for large $m$. To mitigate this, we apply NGMRES($m$) at periodic intervals within the fixed-point iteration. We propose the following alternating NGMRES with depth $m$ and periodicity $p$, denoted by aNGMRES($m,p$) in Algorithm 3. The main idea is that every $p-1$ steps of fixed-point iterations are followed by one step of NGMRES($m$). By doing this, we do not need to solve the least-squares problem at every iteration. We notice that same idea has been applied to alternating Anderson acceleration, see [14, 9, 20, 12]. There are two advantages of this new algorithm. Firstly, aNGMRES($m,p$) solves a least squares problem only at periodic intervals of length $p$, which reduces computational complexity and increases locality. Secondly, it may enhance the robustness against stagnations, which is observed in NGMRES($m$) (see [6]), by making their occurrence less likely. We will explore connections between aNGMRES, GMRES and restarted GMRES (GMRES($m$)) where GMRES is restarted once the Krylov subspace reaches dimension $m$ and the current approximate solution becomes the new initial guess for the next $m$ iterations.
@@
Note that when $p=1$, aNGMRES($m,p$) is NGMRES($m$). When $m=\infty$, then (18) is NGMRES($k-1$), which uses all previous iterates to compute $u_{k}$. We comment that in Algorithm 3, the fixed-point iteration operator $q$ can be any iterative method. However, in this work, we limit ourselves to the Richardson iteration (2) for the purpose of convergence analysis.
@@ 3.1 Comparison between aNGMRES, GMRES, and restarted GMRES
To study the relationship between aNGMRES($m,p$), GMRES, and restarted GMRES, we first introduce some terminologies. We denote the Krylov subspace associated with $A$ and initial residual $r_{0}$ by
@@
Following [13], we define the stagnation index of full GMRES as
@@
The degrade of $r_{0}\neq 0$ with respect to $A$ is defined as (see [15])
@@
We proceed the proof by induction. Note that
@@
For $k=1$, from (19), the equation (21) holds. Assume that for $k\leq s$, $c_{k,k-1}=(-1)^{k}$. Then
@@
From (21), we know the above system has a solution. It follows that $A^{s}r_{0}\in\Delta$. We have $\Delta=\mathcal{K}_{s+1}(A,r_{0})$, which is the desired result. ∎
If we consider aNGMRES($m,p$) with finite $m$ and $p=m+1$ applied to $q(u)=Mu+b$, then aNGMRES($m,m+1$) and restarted GMRES($m+1$) are identical at the end of each periodic interval of length $m+1$. We state it in the following theorem.
@@
$\min_{{\bf\tau}\in\mathbb{R}^{p}}\|Au_{p}-b\|_{2}=\min_{{\bf\tau}\in\mathbb{R}^{p}}\|A(u_{0}+\sum_{i=1}^{p}\tau^{(p-1)}_{i}w_{i})-b\|_{2}=\min_{w\in\mathcal{K}_{p}(A,r_{0})}\|A(u_{0}+w)-b\|_{2},$
@@
The above theorem suggests that aNGMRES($m,m+1$) can be treated as an alternative to restarted GMRES. Note that when $m=0,p=1$, aNGMRES($m,p$) is NGMRES(0), which is the same as GMRES(1) shown in [6]. In Theorem 3.1, all iterates are identical.
@@
Consider aNGMRES($\infty,p$) presented in Algorithm 3 applied to $q(u)=Mu+b$. Assume aNGMRES($\infty,p$) and full GMRES use the same initial guess $u_{0}$ and denote the corresponding sequences generated by them as $\left\{u_{k}\right\}$ and $\left\{u_{k}^{G}\right\}$, respectively. Then,
@@
Moreover, if the stagnation does not occur, then there exists an $s\in\mathbb{N}$ such that
@@
Note that the first $p$ iterates, i.e., $\{u_{j}\}_{j=1}^{p}$, of aNGMRES($\infty,p$) are the same as these of aNGMRES($p-1,p$). From Theorem 3.1, we have
@@
Since $p<\nu(A,r_{0})$ and $p<\eta^{G}$, we have $\mathcal{K}_{p}^{\alpha}=\mathcal{K}_{p}(A,r_{0})$. We proceed the proof by induction. Assume that for $j\leq j_{*}$ the $(j-1)p$th iteration satisfies
@@
After the $(j-1)p$th iteration, we perform $p-1$ successive Richardson iterations. For $i=1,2,\cdots,p-1$, define $u_{(j-1)p+i}=u_{0}+w_{(j-1)p+i}$. Then, for $i=1$
@@
$\min_{{\bf\tau}\in\mathbb{R}^{jp}}\|Au_{jp}-b\|_{2}=\min_{{\bf\tau}\in\mathbb{R}^{jp}}\|A(u_{0}+\sum_{i=1}^{jp}\tau^{(jp-1)}_{i}z_{i})-b\|_{2}=\min_{z\in\mathcal{\bar{K}}}\|A(u_{0}+z)-b\|_{2}.$ (25)
@@
If $(s-1)p<\nu(A,r_{0})<sp$, then GMRES converges to the exact solution at $\nu$th iteration, and
@@
After the $(s-1)p$th iteration, aNGMRES($\infty,p$) processes $(p-1)$th Richardson iteration, and $\mathcal{\bar{K}}_{sp}$ has the same dimension as $\mathcal{K}_{\nu}(A,r_{0})$. Similar to (25), the $(sp)$th iteration of aNGMRES($\infty,p$) is given by
$u_{sp}=q(u_{sp-1})+\sum_{i=0}^{sp-1}\beta^{(sp-1)}_{i}\left(q(u_{sp-1})-u_{sp-1-i})\right),$
where the coefficients, $\beta^{(sp-1)}_{i}$, solve
$\min_{\beta^{(sp-1)}_{i}}\|Au_{sp}-b\|_{2}=\min_{z\in\mathcal{\bar{K}}_{sp}}\|A(u_{0}+z)-b\|_{2}.$
@@
We note that when $p=1$, aNGMRES($\infty$,1) reduces to NGMRES($\infty$). As shown in [6], NGMRES($\infty$) is identical to GMRES if the norm of the residuals of GMRES is strictly decreasing. For $p=1$, our result in (22) is consistent with the findings in [6]. In [9] the author proposed Anderson-type acceleration of Richardson’s iteration (AAR) and showed that the $(jp)$th iteration by full AAR ($m=\infty$) after the Anderson mixing is the same as $(jp)$th GMRES iteration. Now, our theorem makes connection with AAR due to the equivalence to GMRES.
When GMRES stagnation occurs, the convergence behavior of aNGMRES becomes complex. We do not explore this direction further from a theoretical perspective. Instead, we examine it numerically.
@@ 3.2 Convergence analysis of aNGMRES
In the following, we discuss the convergence of aNGMRES. We first consider the fixed-point iteration is contractive.
@@
If $u_{k}$ is obtained from the fixed-point iteration, then $r_{k}=Mr_{k-1}$, and $\|r_{k}\|\leq c\|r_{k-1}\|$. If $u_{k}$ is obtained from one step NGMRES($m$), then from (4), we have $\|r_{k}\|\leq\|r(q(u_{k-1}))\|=\|Mr_{k-1}\|\leq c\|r_{k-1}\|$. Thus, it holds that $\|r_{k}\|\leq c\|r_{k-1}\|$ for every $k$. Thus $\|r_{k}\|$ converges to zero. ∎
The above result emphasizes on the convergence, but does not show by much that aNGMRES can improve the convergence speed of fixed-point iteration. In the following, we derive convergence bounds for some cases. In fact, we observe that even when $M$ is not contractive, aNGMRES may converge.
@@
Let $q(u)=Mu+b$. Consider aNGMRES($m,p$) with $m<p-1$ presented in Algorithm 3. When $M$ is diagonalizable, we have the following error estimate
@@
If $p=m+1$, when $M$ is diagonalizable, we have the following error estimate
@@
When $A$ is SPD, we have the following error estimate
@@
These results can be derived directly from (12), (13) and (17). The second inequality in (27) can be derived easily, see [15]. Thus, we omit it. ∎
When $M$ is diagonalizable and $m<p-1$, if $\epsilon(a,b,m+1)\|M\|^{p}\kappa_{2}(U)<1$, then the subsequence $\left\{r_{jp}\right\}$ of $\left\{r_{k}\right\}$ converges to zero. If $\|M\|<1$, then either $\big{(}\epsilon(a,b,m+1)\kappa_{2}(U)\big{)}^{j}$ or $\big{(}\epsilon(a,b,m+1)\big{)}^{j}$ is the gain obtained by aNGMRES at $jp$ iteration compared with fixed-point iteration.
When $M$ is diagonalizable and $p=m+1$, if $\chi(m+1)\kappa_{2}(U)<1$, then the subsequence $\left\{r_{j(m+1)}\right\}$ of $\left\{r_{k}\right\}$ converges to zero. When $A$ is SPD, we can choose $m$ sufficiently large such that $12\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{m+1}<$ 1, then the subsequence $\left\{r_{j(m+1)}\right\}$ of $\left\{r_{k}\right\}$ converges to zero.

### v2

@@ 2.1.2 Symmetric and positive definite case
In the following, we consider $M$ to be symmetric and positive definite (SPD), i.e., there exists a unitary matrix $U$ such that $M=U\Lambda U^{T}$, where $\Lambda$ is a diagonal matrix. Then, $e_{j+1}=M^{j+1}e_{0}$ and $e_{j+1}-e_{j}=M^{j}(M-I)e_{0}=U\Lambda^{j}(\Lambda-I)U^{T}e_{0}$.
@@
Assume that $M$ is SPD, i.e., $M=U\Lambda U^{T}$, where $U$ is a unitary matrix and $\Lambda$ is a diagonal matrix, the eigenvalues of $M$ are contained in an interval $a_{1},a_{2}]$ that does not include 0 and 1, and $m<k$. Then, the 2-norm of the error $e_{k+1}$ obtained from one-step NGMRES($m$) satisfies
$\|\Theta^{-1}U^{T}e_{k+1}\|\leq\epsilon(a_{1},a_{2},m+1)\cdot\|\Theta^{-1}U^{T}Me_{k}\|,$ (18)
@@
$\leq\min_{p_{m}\in P_{m}}\max_{t\in[a_{1},a_{2}]}\left|1-p_{m}(t)t^{-m-1}(1-t)\right|\|\Theta^{-1}U^{T}Me_{k}\|$
$\leq\min_{p_{m}\in P_{m}}\max_{t\in[1/a_{2},1/a_{1}]}\left|1-p_{m}(1/t)t^{m}(t-1)\right|\|\Theta^{-1}U^{T}Me_{k}\|$
$\leq\min_{p_{m}\in P_{m}}\max_{t\in[1/a_{2},1/a_{1}]}\left|1-p_{m}(t)(t-1)\right|\|\Theta^{-1}U^{T}Me_{k}\|$
$\leq\min_{q_{m+1}\in P_{m+1},q_{m+1}(1)=1}\max_{t\in[1/a_{2},1/a_{1}]}\left|q_{m+1}(t)\right|\|\Theta^{-1}U^{T}Me_{k}\|.$
@@
The second result can be directly obtained from Theorem 2.1 using the fact that $\kappa_{2}(U)=1$. The last result, (20), is obtained from Theorem 2.2.
Note that if the $(k+1)$th iterate is obtained by the fixed-point iteration, then the corresponding residual is $Mr_{k}$. Therefore, the factor $\epsilon(a_{1},a_{2},m+1)$ in (19) provides a quantitative measurement of how much the fixed-point iteration can be improved.
@@
Note that one-step Anderson acceleration for the symmetric case has been studied [26]. The Chebyshev polynomial $\epsilon(a_{1},a_{2},m+1)$ for NGMRES($m$) is of degree $m+1$, but it is $m$ for AA($m$), and $\Theta$ is $\Lambda(\Lambda-I)^{-1}$.
Based on the one-step NGMRES($m$) described above, we next propose a new variant of NGMRES and conduct its convergence analysis.
@@ 3 Alternating NGMRES
Note that NGMRES($m$) requires solving a least-squares problem at each iteration, which is computationally expensive for large $m$. To mitigate this, we apply NGMRES($m$) at periodic intervals within the fixed-point iteration. We propose the following alternating NGMRES with depth $m$ and periodicity $p$, denoted by aNGMRES($m,p$) in Algorithm 3. The main idea is that every p1p-1 steps of fixed-point iterations are followed by one step of NGMRES($m$). By doing this, we do not need to solve the least-squares problem at every iteration. In Algorithm 3, $u_{k}$ defined in (21) is updated using (7). We notice that the same idea has been applied to alternating Anderson acceleration [20, 15, 23, 18]. There are two advantages of this new algorithm. Firstly, aNGMRES($m,p$) solves a least squares problem only at periodic intervals of length $p$, which reduces computational complexity and increases locality. Secondly, it may enhance the robustness against stagnations, which is observed in NGMRES($m$) [10], by making their occurrence less likely. We will explore connections between aNGMRES, GMRES and restarted GMRES (GMRES($m$)) where GMRES is restarted once the Krylov subspace reaches dimension $m$ and the current approximate solution becomes the new initial guess for the next $m$ iterations.
@@
Note that when $p=1$, aNGMRES($m,p$) is NGMRES($m$). When $m=\infty$, then (21) is NGMRES(k1k-1), which uses all previous iterates to compute $u_{k}$. We comment that in Algorithm 3, the fixed-point iteration operator $q$ can be any iterative method. However, in this work, we limit ourselves to the Richardson iteration (2) for the purpose of convergence analysis.
@@ 3.1 Comparison between aNGMRES, GMRES, and restarted GMRES
To study the relationship between aNGMRES($m,p$), GMRES, and restarted GMRES, we first introduce some terminologies. We denote the Krylov subspace associated with AA and the initial residual $r_{0}$ by
@@
Following the work of Potra and Engler [19], we define the stagnation index of full GMRES as follows.
@@
The grade of $r_{0}\neq 0$ with respect to $A$ is defined as [19]
@@
We proceed with the proof by induction. Note that
@@
For $k=1$, from (22), equation (24) holds. Assume that for $k\leq s$, $c_{k,k-1}=(-1)^{k}$. Then
@@
From (24), we know that the above system has a solution. It follows that $A^{s}r_{0}\in\Delta$. We have $\Delta=\mathcal{K}_{s+1}(A,r_{0})$, which is the desired result. ∎
Lemma 3.1 establishes a connection between Krylov subspace of GMRES and the space associated with fixed-point iterations, which is linked to the periodic iterates of aNGMRES. If we consider aNGMRES($m,p$) with finite $m$ and $p=m+1$ applied to $q(u)=Mu+b$, then aNGMRES($m,m+1$) and the restarted GMRES (GMRES($m+1$)) are identical at the end of each periodic interval of length $m+1$. We state it in the following theorem.
@@
The key idea of the proof is to examine the Krylov subspace associated with the aNGMRES iterates and to use the property that the $k$th NGMRES iterate minimizes the residual norm.
@@
$\min_{{\bf\tau}\in\mathbb{R}^{p}}\|Au_{p}-b\|=\min_{{\bf\tau}\in\mathbb{R}^{p}}\|A(u_{0}+\sum_{i=1}^{p}\tau^{(p-1)}_{i}w_{i})-b\|=\min_{w\in\mathcal{K}_{p}(A,r_{0})}\|A(u_{0}+w)-b\|,$
@@
The above theorem suggests that aNGMRES($m,m+1$) can be treated as an alternative to restarted GMRES. Note that when $m=0,p=1$, aNGMRES($m,p$) is NGMRES(0), which is the same as GMRES(1) [10], i.e., $u_{k}=u_{k}^{G(1)}$ for all $k$, as stated in Theorem 3.1.
@@
Consider aNGMRES($\infty,p$) presented in Algorithm 3 applied to $q(u)=Mu+b$. Assume that aNGMRES($\infty,p$) and full GMRES use the same initial guess $u_{0}$ and denote the corresponding sequences generated by them as $\left\{u_{k}\right\}$ and $\left\{u_{k}^{G}\right\}$, respectively. Then,
@@
Moreover, if the stagnation of GMRES does not occur, then there exists a $s\in\mathbb{N}$ such that
@@
Note that the first $p$ iterates, i.e., $\{u_{j}\}_{j=1}^{p}$, of aNGMRES($\infty,p$) are the same as these of aNGMRES(p1,pp-1,p). From Theorem 3.1, we have
@@
Since $p<\nu(A,r_{0})$ and $p<\eta^{G}$, we have $\mathcal{K}_{p}^{\alpha}=\mathcal{K}_{p}(A,r_{0})$. We proceed with the proof by induction. Assume that for $j\leq j_{*}$ the (j1)p(j-1)pth iteration satisfies
@@
After the (j1)p(j-1)pth iteration, we perform p1p-1 successive Richardson iterations. For $i=1,2,\cdots,p-1$, define $u_{(j-1)p+i}=u_{0}+w_{(j-1)p+i}$. Then, for $i=1$
@@
$\min_{{\bf\tau}\in\mathbb{R}^{jp}}\|Au_{jp}-b\|=\min_{{\bf\tau}\in\mathbb{R}^{jp}}\|A(u_{0}+\sum_{i=1}^{jp}\tau^{(jp-1)}_{i}z_{i})-b\|=\min_{z\in\mathcal{\bar{K}}}\|A(u_{0}+z)-b\|.$ (28)
@@
[…]
@@
After the (s1)p(s-1)pth iteration, aNGMRES($\infty,p$) processes (p1)(p-1) Richardson iterations, and $\mathcal{\bar{K}}_{sp}$ has the same dimension as $\mathcal{K}_{\nu}(A,r_{0})$. Similarly to (28), the $(sp)$th iteration of aNGMRES($\infty,p$) is given by
$u_{sp}=q(u_{sp-1})+\sum_{i=0}^{sp-1}\beta^{(sp-1)}_{i}\left(q(u_{sp-1})-u_{sp-1-i})\right).$
Let $\bm{\beta}^{(sp-1)}=[\beta^{(sp-1)}_{0},\beta^{(sp-1)}_{1},\cdots,\beta^{(sp-1)}_{sp-1}]^{T}$, which solves
$\min_{\bm{\beta}^{(sp-1)}}\|Au_{sp}-b\|=\min_{z\in\mathcal{\bar{K}}_{sp}}\|A(u_{0}+z)-b\|.$
@@
We note that when $p=1$, aNGMRES($\infty$,1) reduces to NGMRES($\infty$). As shown in our recent work [10], NGMRES($\infty$) is identical to GMRES if the norm of the residuals of GMRES is strictly decreasing. For $p=1$, our result in (25) is consistent with recent findings [10]. Recently, Lupo Pasini [15] proposed Anderson-type acceleration of Richardson’s iteration (AAR) and showed that the $(jp)$th iteration by full AAR ($m=\infty$) after the Anderson mixing is the same as the $(jp)$th GMRES iteration. Now, our theorem makes connection with AAR due to the equivalence to GMRES.
When GMRES stagnation occurs, the convergence behavior of aNGMRES becomes complex. We do not explore this direction further from a theoretical perspective. Instead, we examine it numerically; see the numerical results of Example 4.2 in subsection 4.2.
@@ 3.2 Convergence analysis of aNGMRES
In the following, we discuss the convergence of aNGMRES. We first consider the fixed-point iteration to be contractive.
@@
If $u_{k}$ is obtained from the fixed-point iteration, then $r_{k}=Mr_{k-1}$, and $\|r_{k}\|\leq c\|r_{k-1}\|$. If $u_{k}$ is obtained from a one-step NGMRES($m$), then from (6), we have $\|r_{k}\|\leq\|r(q(u_{k-1}))\|=\|Mr_{k-1}\|\leq c\|r_{k-1}\|$. Thus, we have $\|r_{k}\|\leq c\|r_{k-1}\|$ for every $k$. Because $c<1$, the sequence $\{\|r_{k}\|\}$ converges to zero. ∎
The above result emphasizes on the convergence, but does not show by how much that aNGMRES can improve the convergence speed of the underlying fixed-point iteration. In the following, we derive convergence bounds for some cases. In fact, we observe that even when $M$ is not contractive, aNGMRES may converge.
@@
Let $q(u)=Mu+b$. Consider aNGMRES($m,p$) with m<p1m<p-1 presented in Algorithm 3. When $M$ is diagonalizable, we have the following error estimate
@@
If $p=m+1$ and $M$ is diagonalizable, we have the following error estimate
@@
If $p=m+1$ and AA is SPD, we have the following error estimate
@@
These results can be derived directly from (15), (16) and (20). The second inequality in (30) can be derived easily by following Section 6.11.3 in the work of Saad [21]. Therefore, we omit it. ∎
When $M$ is diagonalizable and m<p1m<p-1, if $\epsilon(a,b,m+1)\|M\|^{p}\kappa_{2}(U)<1$, then the subsequence $\left\{r_{jp}\right\}$ of $\left\{r_{k}\right\}$ converges to zero. When $M$ is SPD, it is possible that the subsequence of aNGMRES can converge provided that $\epsilon(a,b,m+1)\|M\|^{p}<1$. If $\|M\|<1$, then either $\big(\epsilon(a,b,m+1)\kappa_{2}(U)\big)^{j}$ or $\big(\epsilon(a,b,m+1)\big)^{j}$ is the gain obtained by aNGMRES at $jp$ iteration compared to the fixed-point iteration.
When $M$ is diagonalizable and $p=m+1$, if $\chi(m+1)\kappa_{2}(U)<1$, then the subsequence $\left\{r_{j(m+1)}\right\}$ of $\left\{r_{k}\right\}$ converges to zero. When $A$ is SPD, we can choose $m$ sufficiently large such that $2\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{m+1}<1$, then the subsequence $\left\{r_{j(m+1)}\right\}$ of $\left\{r_{k}\right\}$ converges to zero. The above theorem suggests that both the condition number of AA and the window size $m$ used in aNGMRES influence the rate of convergence of aNGMRES. In particular, a smaller condition number $\kappa$ is associated with improved algorithmic performance.
Remark 3.1.
As previously noted, aNGMRES can be applied to general fixed-point iterations. In this work, however, we restrict our convergence analysis to the specific form q(x)=x+(bAx)q(x)=x+(b-Ax). It would be interesting to extend this analysis to the more general case q(x)=x+P(bAx)q(x)=x+P(b-Ax), where $P$ serves as a preconditioner—leading to a preconditioned version of aNGMRES. A natural question that arises is the relationship between preconditioned aNGMRES and preconditioned GMRES algorithm, including its restarted variants. However, this investigation is beyond the scope of the present work.